# Denoising Complex Covariance Matrices with Hybrid ResNet and Random Matrix Theory: Cryptocurrency Portfolio Applications

### v1

@@ Abstract
Covariance matrices estimated from short, noisy, and non-Gaussian financial time series—particularly cryptocurrencies—are notoriously unstable. Empirical evidence indicates that these covariance structures often exhibit power-law scaling, reflecting complex and hierarchical interactions among assets. Building on this insight, we propose a power-law covariance model to characterize the collective dynamics of cryptocurrencies and develop a hybrid estimator that integrates Random Matrix Theory (RMT) with Residual Neural Networks (ResNets). The RMT component regularizes the eigenvalue spectrum under high-dimensional noise, while the ResNet learns data-driven corrections to recover latent structural dependencies. Monte Carlo simulations show that ResNet-based estimators consistently minimize both Frobenius and minimum-variance (MV) losses across diverse covariance models. Empirical experiments on 89 cryptocurrencies (2020–2025), using a training period ending at the local BTC maximum in November 2021 and testing through the subsequent bear market, demonstrate that a two-step estimator combining hierarchical filtering with ResNet corrections yields the most profitable and balanced portfolios, remaining robust under market regime shifts. These findings highlight the potential of combining RMT, deep learning, and power-law modeling to capture the intrinsic complexity of financial systems and enhance portfolio optimization under realistic conditions.
@@ 1 Introduction
Random Matrix Theory (RMT) has played a pivotal role in elucidating complex interactions in nuclear physics [1], revealing quantum chaos signatures [2], and describing quantum transport phenomena [3]. More recently, RMT has been successfully applied to denoise empirical covariance matrices and enhance optimal asset allocation in portfolio theory [4, 5, 6, 7]. Within this framework, several methodologies have been proposed to identify meaningful signals and suppress noise originating from the finite size of empirical samples [4, 8]. Furthermore, renewed attention has arisen in mathematical statistics, where advanced techniques grounded in free probability and deterministic equivalents have provided rigorous analytical tools for high-dimensional covariance estimation [9, 6, 10].
These approaches share the common characteristic of focusing primarily on the eigenvalues of the covariance matrix, whose empirical distribution asymptotically converges to a deterministic function. In contrast, the role of eigenvectors has been largely secondary within these frameworks. In essence, RMT-based estimators improve covariance estimation by applying nonlinear shrinkage transformations to the eigenvalues while keeping the eigenvectors fixed [11, 5, 12]. However, it has been demonstrated that the distance between the leading sample eigenvectors and their true population counterparts diverges without bound in high-dimensional settings [13].
@@
Therefore, this work proposes to exploit information contained in the eigenvectors to improve covariance matrix estimation. Our approach performs a spectral decomposition of the covariance matrix, estimating the eigenvectors through a state-of-the-art deep neural network model, while the eigenvalues are obtained via the nonlinear shrinkage method based on Random Matrix Theory (RMT) [7]. In particular, we employ the Residual Neural Network (ResNet) architecture [14, 15], which is well-suited for deep architectures as it mitigates computational challenges by introducing shortcut connections that facilitate efficient gradient propagation toward optimal solutions.
@@
This proposal is benchmarked against two alternative covariance estimators: one entirely based on the ResNet architecture and another fully derived from Random Matrix Theory (RMT). Additionally, we include the two-step estimator introduced in [16], which incorporates the nested hierarchical organization of financial markets into advanced covariance estimation techniques. Since the main goal of this study is to enhance covariance matrix estimation for financial systems, we first introduce models designed to reproduce the stylized facts of market interactions. To the best of our knowledge, this is the first work to propose a power-law covariance model whose scree plot of eigenvalues follows a scaling behavior, thereby capturing the complex dynamics of financial markets. For comparison, we also consider a fully hierarchical model [17] and a block-diagonal model [16].
To validate the models in practice, we have considered a dataset composed of financial instruments known as cryptocurrencies. These assets are characterized by complex features that push traditional statistical methods to their limits. Specifically, cryptocurrencies do not follow a Gaussian distribution, exhibit heavy tails, abrupt jumps, asymmetry, and overall complex dynamics [18].
@@
Here, we carefully excluded pseudo-cryptocurrencies—i.e., coins that are merely replicas of fiat money—and retained only genuine blockchain projects or those that replicate exchange rates. Additionally, we pushed the estimators, particularly the deep learning ones, to their limits by training on a bull market period and testing on a bear market. This setup allows us to evaluate the model’s ability to handle structural changes and market regime shifts.
We find that both the nested hierarchy and power-law models adequately characterize the complex interactions within cryptocurrency markets. Moreover, both in simulations and in empirical cryptocurrency data, hybrid models—and more generally, deep learning-based models—significantly improve the estimation of the population covariance matrix on one hand, and, on the other, enhance financial metrics in out-of-sample walk-forward analyses.
Although the covariance matrix is not the only element involved in an investment strategy, its accurate estimation enables better control of volatility under different scenarios—not only within portfolio theory. For instance, one of the most recent and prestigious forecasting competitions found that the main determinant of predictive accuracy was the ability to control volatility, regardless of model complexity [19].
On the other hand, while recent studies have explored the use of deep learning for estimating eigenvalues within invariant estimator frameworks [20], or through reinforcement learning approaches to shrinkage estimation [21], to the best of our knowledge, this is the first work to propose a hybrid method that integrates the theoretical robustness of RMT with the representational power of deep neural networks applied to eigenvectors, specifically leveraging the ResNet architecture.
@@ 2.2 Minimum Variance Portfolio (MVP)
The mean–variance allocation strategy proposed by Markowitz [22] seeks to solve the following quadratic optimization problem in order to minimize portfolio risk for a given level of expected return [23]:
@@
The reduced problem minimizes portfolio volatility and is referred to as the Minimum Variance Portfolio (MVP). Moreover, it is possible to incorporate a no short-selling constraint by formulating the standard quadratic programming (QP) problem. In this case, the portfolio weights satisfy $\mathbf{w}\geq 0$, and a numerical solution is required [24]. This constrained portfolio is also known as the long-only MVP, as it does not permit short positions or negative weights. In the following analysis, we will denote this portfolio as MVP+.
@@ 3.1 Random matrix denoising
which is an unbiased estimator of $\boldsymbol{\Sigma}$ when the number of variables $p$ is fixed and the number of observations $n\to\infty$ [25].
A non-linear shrinkage formula that minimizes the Frobenius loss has been proposed by Ledoit and Péché [5], based on principles from RMT [26]. In high-dimensional settings, the covariance matrix is estimated as
@@ 3.2 Residual Neural Network denoising
An estimator based on machine learning is proposed, employing the Residual Neural Network (ResNet) architecture [14]. The implementation builds upon Convolutional Neural Networks (CNNs) [27], which are designed to process grid-structured inputs and exploit strong spatial dependencies within local regions of the grid. ResNet improves CNNs by introducing a novel neural architecture that incorporates skip connections [28], allowing information from one layer to propagate to non-contiguous layers. Specifically, ResNet contains connections between layer $i$ and layer $(i+r)$ for $r>1$, where $i$ is an arbitrary layer and $r$ denotes the skip length. Figure 1(a) illustrates the architecture of the residual block used for the covariance matrix denoising, setting $r=2$.
@@
The CNN learning process operates as follows: a filter is applied to the input layer to detect patterns, producing a feature map as the output. An activation function is then applied element-wise after each layer to introduce nonlinearity, regulate signal propagation, and constrain the output within a desired range of values.
In the residual module, the input layer consists of 64 filters of size $3\times 3$, producing 64 feature maps with rectified linear unit (ReLU) nonlinearity. This base representation is then refined by two subsequent stacked 2D convolutional layers (Conv2D) to extract deeper features. The first Conv2D layer employs a ReLU activation function, whereas the second uses a linear activation function. The add block implements the skip connection, combining the input (residual) with the output of the two Conv2D layers. In this way, the network learns only the residual correction rather than the full mapping, mitigating the vanishing gradient problem. The practical purpose of the skip connections is to enable effective gradient flow, allowing the learning algorithm to adaptively choose the level of nonlinearity applied to a given input. Finally, the last block of the residual module applies a ReLU activation. The complete network architecture, shown in Figure 1(b), is constructed by repeating this loop and stacking ten basic residual blocks. Using this ResNet architecture, and applying appropriate symmetrization along with a positive semidefinite (PSD) transformation, it becomes possible to learn and remove the noise present in the empirical covariance matrix. We denote the resulting estimator as $\boldsymbol{\Xi}^{\text{CNN}}$, reflecting the CNN-based structure of its inner layers.
@@
$\boldsymbol{\Xi}^{H}=\boldsymbol{\Xi}^{\text{CNN}}(\mathbf{V})\,\boldsymbol{\Xi}(\mathbf{\Lambda})\,\boldsymbol{\Xi}^{\text{CNN}}(\mathbf{V})^{\top},$ (13)
where $\boldsymbol{\Xi}(\mathbf{V})$ and $\boldsymbol{\Xi}(\mathbf{\Lambda})$ denote the application of the operator $\boldsymbol{\Xi}$ to the matrix of eigenvectors and the diagonal matrix of eigenvalues, respectively. This hybrid approach combines the data-driven flexibility of ResNet to capture complex dependencies in the eigenvectors, while exploiting the theoretical guarantees of RMT for consistent eigenvalue estimation in high-dimensional settings.
@@ 3.3 Hierarchical clustering denoising
A novel approach to covariance matrix estimation was proposed by Tumminello et al. [29] using a hierarchical clustering algorithm. The procedure begins by transforming the empirical covariance matrix $\mathbf{S}$ into the corresponding correlation matrix $\mathbf{C}$. Next, the transformation
@@
is applied, which satisfies the axioms of a distance measure, where $\mathbf{1}$ is a $p$-dimensional vector of ones. A dendrogram is then constructed from $\mathbf{D}$ using Average Linkage Clustering Analysis (ALCA) [25], and the distance $\rho$ between clusters at each hierarchical level is computed. This procedure yields a dissimilarity matrix $\mathbf{D}(\rho)$ as a function of $\rho$, from which the filtered covariance matrix is recovered via the inverse transformation:
@@
A state-of-the-art estimator designed to address both the heterogeneous structure of financial markets and the challenges of high-dimensional settings was proposed in [16]. The core idea is to first apply our proposed single-step estimators, followed by a hierarchical filtering step. We consider the following combinations:
@@
The mathematical justification for this two-step estimator is discussed in [16] and is based on the following arguments: (i) applying an RMT-based estimator in the first step reduces the noise and estimation error associated with the largest eigenvalues, and (ii) applying a subsequent filtering step via hierarchical clustering mitigates the inconsistency of the top eigenvectors, which is inherently present in RMT covariance estimators [13]. Here, this reasoning is extended to the newly proposed estimators $\boldsymbol{\Xi}^{\text{CNN}}$ and $\boldsymbol{\Xi}^{H}$, where the second step can be interpreted as a regularization procedure to avoid overfitting.

### v2

@@ Abstract
Covariance matrices estimated from short, noisy, and non-Gaussian financial time series are notoriously unstable. Empirical evidence suggests that such covariance structures often exhibit power-law scaling, reflecting complex, hierarchical interactions among assets. Motivated by this observation, we introduce a power-law covariance model to characterize collective market dynamics and propose a hybrid estimator that integrates Random Matrix Theory (RMT) with deep Residual Neural Networks (ResNets). The RMT component regularizes the eigenvalue spectrum in high-dimensional noisy settings, while the ResNet learns data-driven corrections that recover latent structural dependencies encoded in the eigenvectors. Monte Carlo simulations show that the proposed ResNet-based estimators consistently minimize both Frobenius and minimum-variance losses across a range of population covariance models. Empirical experiments on 89 cryptocurrencies over the period 2020–2025, using a training window ending at the local Bitcoin peak in November 2021 and testing through the subsequent bear market, demonstrate that a two-step estimator combining hierarchical filtering with ResNet corrections produces the most profitable and well-balanced portfolios, remaining robust across market regime shifts. Beyond finance, the proposed hybrid framework applies broadly to high-dimensional systems described by low-rank deformations of Wishart ensembles, where incorporating eigenvector information enables the detection of multiscale and hierarchical structure that is inaccessible to purely eigenvalue-based methods.
@@ 1 Introduction
Random Matrix Theory (RMT) has played a pivotal role in elucidating complex interactions in nuclear physics [mehta2004random], revealing signatures of quantum chaos [bohigas1984characterization], and describing quantum transport phenomena [beenakker1997random]. The Wishart distribution describes the sample covariance matrix of multivariate Gaussian data [wishart1928generalised]. Within the framework of RMT, such covariance matrices are commonly referred to as members of the Wishart Orthogonal Ensemble (WOE). The Wishart distribution can be viewed as a natural multivariate generalization of the relationship between the univariate normal distribution and the chi-square distribution [pena2002analisis].
Wishart-type random matrix models naturally arise in a wide range of physical and applied contexts, including antenna selection in communication technologies [sadek2007active], nuclear physics [fyodorov1997statistics], quantum chromodynamics [verbaarschot1994spectrum], and nonintersecting Brownian motions [schehr2008exact].
@@
More recently, RMT has been successfully applied to the denoising of empirical covariance matrices of Wishart type, leading to improved optimal asset allocation within portfolio theory [laloux1999noise, ledoit2011eigenvectors, bun2017cleaning, ledoit2020analytical]. Within this framework, several methodologies have been proposed to identify meaningful signals and suppress noise arising from the finite size of empirical samples [laloux1999noise, potters2005financial]. Furthermore, renewed attention has emerged in mathematical statistics, where advanced techniques grounded in free probability theory and deterministic equivalents have provided rigorous analytical tools for high-dimensional covariance estimation [yang2015robust, bun2017cleaning, Burda2022].
In essence, RMT-based estimators improve covariance estimation by applying nonlinear shrinkage transformations to the eigenvalue spectrum while keeping the eigenvectors fixed [ledoit2004well, ledoit2011eigenvectors, ledoit2022power]. These estimators are known as Rotationally Invariant Estimators (RIEs). From this perspective, RMT-based estimators exploit the asymptotic properties of the eigenvalue distribution of the sample covariance matrix, which converges to a deterministic limit in the high-dimensional regime $N/T=\mathcal{O}(1)$ [bun2017cleaning, Burda2022], corresponding to a system of $N$ time series observed over $T$ periods.
Complementary to this spectral approach, the effectiveness of filtering correlation matrices via hierarchical clustering has been extensively documented, particularly in the context of portfolio optimization in finance [tumminello2007hierarchically, tumminello2007spectral, tola2008cluster, pantaleo2011improved, bongiorno2021covariance, bongiorno2022reactive, garcia2023two, garcia2024random, garcia2025high]. At the same time, recent advances in machine learning have opened new avenues for high-dimensional covariance estimation, including deep learning–based methods for eigenvalue regularization within invariant estimator frameworks [bongiorno2025end] and reinforcement learning approaches to adaptive shrinkage estimation [mattera2023shrinkage].
Despite these advances, the role of eigenvectors has remained largely secondary within most existing frameworks. In particular, many denoising techniques rely on unstructured assumptions about the population covariance matrix that fail to reproduce the stylized facts of complex systems, thereby substantially limiting their range of applicability. When structure is introduced, as in the works of [garcia2023two, garcia2025high], both nested hierarchical covariance models and block-diagonal covariance structures give rise to eigenvalue spectra consistent with the so-called spiked covariance matrix model [johnstone2001distribution, donoho2018optimal], namely covariance matrices characterized by a small number of large, isolated eigenvalues that are well separated from the bulk of the spectrum.
A key feature of spiked covariance models is top eigenvector inconsistency [donoho2018optimal], meaning that the angle between sample eigenvectors and their corresponding population eigenvectors converges to a non-zero limit. This implies that filtering procedures based on Rotationally Invariant Estimators (RIEs), which retain sample eigenvectors, may fail to capture important structural properties of the population covariance matrix.
A further limitation of RIE methods arises for sample eigenvectors associated with small eigenvalues: these eigenvectors typically have components spread across the entire set of variables, producing orientations that differ substantially from the localized eigenvectors linked to the correlated dynamics of smaller, tightly connected groups of elements.
@@
The above observations motivate an alternative filtering procedure for spiked correlation matrices that combines machine learning with hierarchical clustering. In particular, we propose to explicitly exploit the information contained in the eigenvectors in order to improve covariance matrix estimation. Our approach performs a spectral decomposition of the covariance matrix, estimating the eigenvectors through a state-of-the-art deep neural network model, while the eigenvalues are obtained via the nonlinear shrinkage method based on Random Matrix Theory (RMT) [ledoit2020analytical]. In particular, we employ the Residual Neural Network (ResNet) architecture [he2016deep, he2020resnet], which is well-suited for deep architectures as it mitigates computational challenges by introducing shortcut connections that facilitate efficient gradient propagation toward optimal solutions.
@@
This proposal is benchmarked against two alternative covariance estimators: one entirely based on the ResNet architecture and another fully derived from Random Matrix Theory (RMT). Additionally, we include the two-step estimator introduced in [garcia2023two], which incorporates the nested hierarchical organization of financial markets into advanced covariance estimation techniques. To the best of our knowledge, this is the first work to propose a hybrid estimation framework that combines the theoretical robustness of Random Matrix Theory with the representational capacity of deep neural networks applied to eigenvector information, specifically leveraging a ResNet architecture.
Since no analytical results are currently available regarding the denoising optimality of the proposed hybrid estimator—nor for the hierarchical clustering estimator or its two-step variant—we assess their performance through a comprehensive set of Monte Carlo experiments under a variety of complex population covariance structures. In particular, we consider a fully hierarchical covariance model [garcia2025high] and a block-diagonal covariance model [garcia2023two]. In addition, we introduce a novel population covariance model designed to reproduce key stylized facts of financial market interactions. This model generates a covariance matrix whose eigenvalue spectrum exhibits power-law scaling behavior in the scree plot, thereby capturing salient features of the multiscale dynamics observed in financial markets.
@@
To validate the models in practice, we have considered a dataset composed of financial instruments known as cryptocurrencies. These assets are characterized by complex features that push traditional statistical methods to their limits. Specifically, cryptocurrencies do not follow a Gaussian distribution, exhibit heavy tails, abrupt jumps, asymmetry, and overall complex dynamics [Watorek2021Multiscale]. Here, we carefully excluded pseudo-cryptocurrencies—i.e., coins that are merely replicas of fiat money—and retained only genuine blockchain projects or those that replicate exchange rates. Additionally, we pushed the estimators, particularly the deep learning ones, to their limits by training on a bull market period and testing on a bear market. This setup allows us to evaluate the model’s ability to handle structural changes and market regime shifts.
We find that both the nested hierarchical and power-law covariance models provide an adequate characterization of the complex interactions observed in cryptocurrency markets. Moreover, in both simulation studies and empirical cryptocurrency data, hybrid approaches—and, more broadly, deep learning–based methods—lead to statistically meaningful improvements in population covariance matrix estimation and, in turn, to enhanced out-of-sample financial performance in walk-forward analyses. Although the covariance matrix is not the sole ingredient of an investment strategy, its accurate estimation plays a central role in controlling portfolio volatility across different market conditions, with implications that extend beyond the classical portfolio optimization paradigm [makridakis2024avoiding].
Nevertheless, it is important to emphasize that the scope of the proposed approach is not restricted to financial systems. More generally, any high-dimensional system whose dependence structure can be modeled as a low-rank deformation of a Wishart ensemble may benefit from hybrid covariance estimators that explicitly incorporate eigenvector information. Such hybrid approaches are particularly relevant for systems that inherently exhibit complex nested, hierarchical, or multiscale structures, where purely eigenvalue-based denoising techniques may fail to capture essential geometric and structural properties encoded in the eigenvectors. From a statistical-physics perspective, this corresponds to exploiting information beyond spectral outliers—namely, the collective organization of eigenmodes—which plays a central role in inference problems exhibiting detectability phase transitions [johnstone2001distribution, bloemendal2013limits, feral2009largest].
@@ 2.2 Minimum Variance Portfolio (MVP)
The mean–variance allocation strategy proposed by Markowitz [markowitz1952] seeks to solve the following quadratic optimization problem in order to minimize portfolio risk for a given level of expected return [roncalli2013introduction]:
@@
The reduced problem minimizes portfolio volatility and is referred to as the Minimum Variance Portfolio (MVP). Moreover, it is possible to incorporate a no short-selling constraint by formulating the standard quadratic programming (QP) problem. In this case, the portfolio weights satisfy $\mathbf{w}\geq 0$, and a numerical solution is required [boyd2004convex]. This constrained portfolio is also known as the long-only MVP, as it does not permit short positions or negative weights. In the following analysis, we will denote this portfolio as MVP+.
@@ 3.1 Random matrix denoising
which is an unbiased estimator of $\boldsymbol{\Sigma}$ when the number of variables $p$ is fixed and the number of observations $n\to\infty$ [johnson2002applied].
A non-linear shrinkage formula that minimizes the Frobenius loss has been proposed by Ledoit and Péché [ledoit2011eigenvectors], based on principles from RMT [potters2020first]. In high-dimensional settings, the covariance matrix is estimated as
@@ 3.2 Residual Neural Network denoising
An estimator based on machine learning is proposed, employing the Residual Neural Network (ResNet) architecture [he2016deep]. Starting from the sample covariance matrix $\mathbf{S}$, the ResNet applies repeated residual updates of the form
$\mathbf{S}\;\mapsto\;\mathbf{S}+\mathcal{F}(\mathbf{S}),$ (12)
where $\mathcal{F}$ denotes a generic neural network module. In the present implementation, $\mathcal{F}$ is constructed using Convolutional Neural Networks (CNNs) [brownlee2018deep], which are well suited for processing grid-structured inputs and for exploiting local spatial dependencies within the input representation. ResNet improves CNNs by introducing a novel neural architecture that incorporates skip connections [nielsen2015neural], allowing information from one layer to propagate to non-contiguous layers. Specifically, ResNet contains connections between layer $i$ and layer $(i+r)$ for $r>1$, where $i$ is an arbitrary layer and $r$ denotes the skip length.
The CNN learning process operates as follows: a filter is applied to the input to detect local patterns, producing a corresponding feature map. An activation function is then applied element-wise after each layer to introduce nonlinearity, regulate signal propagation, and constrain the output to a desired range of values. Figure 1(a) illustrates the architecture of the residual block used for the covariance matrix denoising, setting $r=2$.
@@
Here, the residual module is constructed with an initial convolutional layer comprising 64 filters of size $3\times 3$, producing 64 feature maps followed by a rectified linear unit (ReLU) activation. This base representation is subsequently refined through two stacked two-dimensional convolutional layers (Conv2D) to extract higher-level features. The first Conv2D layer employs a ReLU activation function, whereas the second uses a linear activation function.
The add block implements the skip connection by combining the module input (the residual) with the output of the stacked Conv2D layers. In this formulation, the network learns only the residual correction rather than the full mapping, thereby mitigating the vanishing gradient problem. More generally, skip connections facilitate effective gradient propagation, allowing the optimization procedure to adaptively regulate the degree of nonlinearity applied to a given input.
Finally, the residual module applies a ReLU activation at its output. The complete network architecture, illustrated in Fig. 1(b), is obtained by stacking ten such residual blocks. Using this ResNet architecture, together with appropriate symmetrization and a positive semidefinite (PSD) projection, the network learns to attenuate noise in the empirical covariance matrix. We denote the resulting estimator by $\boldsymbol{\Xi}^{\mathrm{CNN}}$, highlighting the CNN-based structure of its internal layers.
@@
$\boldsymbol{\Xi}^{H}=\boldsymbol{\Xi}^{\text{CNN}}(\mathbf{V})\,\boldsymbol{\Xi}^{\text{LP}}(\mathbf{\Lambda})\,\boldsymbol{\Xi}^{\text{CNN}}(\mathbf{V}^{\top}),$ (14)
where $\boldsymbol{\Xi}^{\text{CNN}}(\mathbf{V})$ and $\boldsymbol{\Xi}^{\text{LP}}(\mathbf{\Lambda})$ denote the application of the operators $\boldsymbol{\Xi}^{\text{CNN}}$ and $\boldsymbol{\Xi}^{\text{LP}}$ to the matrix of eigenvectors and the diagonal matrix of eigenvalues, respectively. This hybrid approach combines the data-driven flexibility of ResNet to capture complex dependencies in the eigenvectors, while exploiting the theoretical guarantees of RMT for consistent eigenvalue estimation in high-dimensional settings.
@@ 3.3 Hierarchical clustering denoising
A novel approach to covariance matrix estimation was proposed by Tumminello et al. [tumminello2007kullback] using a hierarchical clustering algorithm. The procedure begins by transforming the empirical covariance matrix $\mathbf{S}$ into the corresponding correlation matrix $\mathbf{C}$. Next, the transformation
@@
is applied, which satisfies the axioms of a distance measure, where $\mathbf{1}$ is a $p$-dimensional vector of ones. A dendrogram is then constructed from $\mathbf{D}$ using Average Linkage Clustering Analysis (ALCA) [johnson2002applied], and the distance $\rho$ between clusters at each hierarchical level is computed. This procedure yields a dissimilarity matrix $\mathbf{D}(\rho)$ as a function of $\rho$, from which the filtered covariance matrix is recovered via the inverse transformation:
@@
A state-of-the-art estimator designed to address both the heterogeneous structure of financial markets and the challenges of high-dimensional settings was proposed in [garcia2023two]. The core idea is to first apply our proposed single-step estimators, followed by a hierarchical filtering step. We consider the following combinations:
@@
The mathematical justification for this two-step estimator is discussed in [garcia2023two] and is based on the following arguments: (i) applying an RMT-based estimator in the first step reduces the noise and estimation error associated with the largest eigenvalues, and (ii) applying a subsequent filtering step via hierarchical clustering mitigates the inconsistency of the top eigenvectors, which is inherently present in RMT covariance estimators [donoho2018optimal]. Here, this reasoning is extended to the newly proposed estimators $\boldsymbol{\Xi}^{\text{CNN}}$ and $\boldsymbol{\Xi}^{H}$, where the second step can be interpreted as a regularization procedure to avoid overfitting.